From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.gov)
Date: Thu Nov 07 2002 - 16:08:24 EST
The HPSS test were concluded yesterday- from our perspective it went ok. Daq syncing ran wihtout a hith 6 Mb/sec over 24 hours and the crs farm was pretty busy. Phobos+phenix were also satisfied, albeit star could only write data at ~ 10 Mb/sec. Hopefulkly there is not a need for more tests. The new hpss sotfware has one hitch - the utility that previuously was used to check free hpss staging disk space (32Gb limit) is not working and needs redoing. In addition we will also have the staging/spool disk increased since rcf has a large pool of leftover disk from the recent upgrades. In this regard - sometimes next week an additional server (rmine002 ?) will be added, and the serving of the data disk split. (data001-006 on one , and the remainin on the other. the changeover should take only about 1/2 hour and can be done once some additional network indfrastructure is in place. /fv ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov
This archive was generated by hypermail 2.1.5 : Thu Nov 07 2002 - 16:19:57 EST